OpenStudy (anonymous):

What is the solution of the system of equations? somebody please help. can someone show me work? 3x+2y+z= 7 5x+5y+4z=3 3x+2y+3z=1

jhonyy9 (jhonyy9):

how you need to solve it with sustitution method or elimination method ?

OpenStudy (anonymous):

I found z by using the equations 1 and 3 and it doesnt tell i need help with x and y

jhonyy9 (jhonyy9):

ok i see it you can get the value of z using first end 3rd equation so how and how many you have got for z ?

OpenStudy (anonymous):

i got z=-3 and i know the answer for y is -1 and x is 4 but i dont know how to get them

jhonyy9 (jhonyy9):

yes z=-3 is right sure so than this value of z=-3 you can using there in first and second equation so than what will get if you us this value of z there ? what will get for first equation and what for second ?

OpenStudy (anonymous):

wait how would set the up?

OpenStudy (anonymous):

this?

jhonyy9 (jhonyy9):

3x+2y-3=7 5x+5y+4*(-3)=3 so can you make the calcule here ?

jhonyy9 (jhonyy9):

courage this is easy

OpenStudy (anonymous):

how do you figure this out? i feel like an idiot right now

jhonyy9 (jhonyy9):

in the first equation assun to both sides 3 and will get how many ? so and for second equation multiplie 4 by (-3) and will get how many ?

jhonyy9 (jhonyy9):

so you dont wann making the calcules here for ?

OpenStudy (anonymous):

-12 for the second equation?

jhonyy9 (jhonyy9):

but how will be firstly the first equation ?

jhonyy9 (jhonyy9):

3x+2y=10 do you understand how i have got this result ?

OpenStudy (anonymous):

oh okay! you added plus three to the 7? and so the 3 crosses out then?

jhonyy9 (jhonyy9):

yes sure so than how you get the second equation ?

OpenStudy (anonymous):

5x+5y-12=-9

jhonyy9 (jhonyy9):

equal 3 and not -9

OpenStudy (anonymous):

okay so what do i do after that?

jhonyy9 (jhonyy9):

so will get 5x+5y-12=3 so now add to both sides 12 and will get ????

OpenStudy (anonymous):

5x+5y=15

jhonyy9 (jhonyy9):

yes so now there are two equation 3x+2y=10 5x+5y=15 yes ?

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

okay! got it so far

jhonyy9 (jhonyy9):

can you solve these system for x and y ?

OpenStudy (anonymous):

yeah let me work it out

OpenStudy (anonymous):

ok now im confused

jhonyy9 (jhonyy9):

3x+2y=10 5x+5y=15 so first of all how do you see what can you factorizing out in the second equation ?

OpenStudy (anonymous):

15?

jhonyy9 (jhonyy9):

why 15 ? and why not just 5 ???

OpenStudy (anonymous):

5

jhonyy9 (jhonyy9):

yes 5 right so than how will be the second equation ?

OpenStudy (anonymous):

im sorry what do you mean?

jhonyy9 (jhonyy9):

5x+5y=15 i have wrote above that you can factorizing out the 5 from the both sides but i mean that you can divide both sides by 5 and so will get 5x+5y=15 /5 x+y=3 ok ?

OpenStudy (anonymous):

okay i see what your ssaying

jhonyy9 (jhonyy9):

so than now there are 3x+2y=10 x+y=3 can you solve it now ?

OpenStudy (anonymous):

what will you solve first the x or y? do you need to get rid of one?

OpenStudy (anonymous):

i got it! i got the answr for y! y=-1

jhonyy9 (jhonyy9):

how ?

jhonyy9 (jhonyy9):

3x+2y=10 x+y=3 --- y=3-x than this substitute in place of y in the first equation 3x+2(3-x)=10 3x+6-2x=10 x+6=10 x=10-6 x=4 ok ?

OpenStudy (anonymous):

ah thats why i just for x! i get it thank you!

jhonyy9 (jhonyy9):

so than 4 +y=3 y=3-4 y=-1 so than result that you have got it right ok ?

OpenStudy (anonymous):

yup! thats what i did, thank you so much

jhonyy9 (jhonyy9):

was my pleasure yw good luck bye